OpenStudy (anonymous):

Write a function rule for the following relation: {(1, 1), (2, 3), (3, 9), (4, 27)}. Graph the function. Write a paragraph describing how you found the function rule for the relation. Include whether or not the function is linear or nonlinear and how you know. If this pattern were to continue, what would

ganeshie8 (ganeshie8):

see any relation between x and y coordinates ?

ganeshie8 (ganeshie8):

when x = 1, y = 1 when x = 2, y = 3 when x = 3, y = 9 when x = 4, y = 27

ganeshie8 (ganeshie8):

??

ganeshie8 (ganeshie8):

look at the y-coordinates, they should look familiar to you hmm ?

OpenStudy (anonymous):

i dont understand ?

ganeshie8 (ganeshie8):

when x = 1, y = 1 = 3^0 when x = 2, y = 3 = 3^1 when x = 3, y = 9 = 3^2 when x = 4, y = 27 = 3^3

ganeshie8 (ganeshie8):

the y-coordinates are simply the powers of 3! thats the pattern ok ?